van der Waals equation:(P+(n^2a/v^2))(v-nb)=nRT
How do I solve for P?

Answer by jorel1380(3719)

(P+(n^2a/v^2)(v-nb)=nRT
(P+(n2a/v2)=nRT/(v-nb)
P=nRT/(v-nb)+n2a/v2..
Who the hell is van der Waals?